OpenStudy (anonymous):

Let f(x)= √3+2x f'(2) ?

OpenStudy (anonymous):

take the derivative i guess

OpenStudy (anonymous):

replace x by 2 compute

OpenStudy (anonymous):

then evaluate it at 2

OpenStudy (anonymous):

oops what lagrange said

OpenStudy (anonymous):

???

OpenStudy (anonymous):

?????????????????????????????????????????????????????????????????????????????????????????????????????

OpenStudy (anonymous):

assuming this is \[f(x)=\sqrt{3+2x}\] then \[f'(x)=\frac{1}{\sqrt{3+2x}}\]

OpenStudy (anonymous):

replace x by 2, see what you get

OpenStudy (anonymous):

Aha i see thanks satellite73